OpenStudy (anonymous):

]Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -1/2x + 4. y = 2x + 8 y = 2x y = -2x - 8 y = -2x

OpenStudy (anonymous):

y -4= 2(x+2) y=2x +8

OpenStudy (anonymous):

LINE GIVEN HAS SLOPE m1=-1/2 line perp have slopem2=-1/m1= 2 so the required line passing through -2,4 having slope m2=2